# Learning to Maximize Mutual Information for Chain-of-Thought Distillation

### v2

@@ Abstract
Knowledge distillation, the technique of transferring knowledge from large, complex models to smaller ones, marks a pivotal step towards efficient AI deployment. Distilling Step-by-Step (DSS), a novel method utilizing chain-of-thought (CoT) distillation, has demonstrated promise by imbuing smaller models with the superior reasoning capabilities of their larger counterparts. In DSS, the distilled model acquires the ability to generate rationales and predict labels concurrently through a multi-task learning framework. However, DSS overlooks the intrinsic relationship between the two training tasks, leading to ineffective integration of CoT knowledge with the task of label prediction. To this end, we investigate the mutual relationship of the two tasks from Information Bottleneck perspective and formulate it as maximizing the mutual information of the representation features of the two tasks. We propose a variational approach to solve this optimization problem using a learning-based method. Our experimental results across four datasets demonstrate that our method outperforms the state-of-the-art DSS. Our findings offer insightful guidance for future research on language model distillation as well as applications involving CoT. Code and models will be released soon.
@@ 3 Methodology
This section starts with preliminaries of IB. Following it, we formulate our CoT distillation idea within the IB and propose a learning approach to optimize MI.
@@ 3.1 Preliminaries
Under the DSS framework, a task prefix [predict] and [explain] will be pretended to the input text text, corresponding to the label prediction and rational generation task, respectively. In the label prediction task, given the input [predict] + text, and the predictive labels $\mathbf{Y}$, a representation feature $\mathbf{V},\mathbf{V}\in\mathbb{R}^{d},$ is trained under $\mathbf{Y}$. In the rationale generation task, given the input [explain] + text and rationale label $\mathbf{R}$, a representation feature $\mathbf{Z},\mathbf{Z}\in\mathbb{R}^{d},$ is trained under $\mathbf{R}$.
Considered the limited CoT capacity of smaller LLMs that are less than 10B parameters, our goal is to distill CoT knowledge of larger LLMs to $Z$ which is also maximally informative to $Y$. In order to achieve this goal, on the basis of IB Tishby and Zaslavsky (2015); Zhang et al. (2022a); Wang et al. (2019), we can model the DSS as following:
@@
where sampling observations $z\thicksim\mathbf{Z}$ and $v\thicksim\mathbf{V}$. $p(\cdot)$ is probability distribution.
To encourage CoT distillation to focus on the information represented in label $\mathbf{Y}$, we propose IB to enforce an upper bound $I_{c}$ to the information flow from the representation features $V$ to the representation features $Z$, achieved by maximizing the following objective:
@@
By using Lagrangian objective, IB allows $Z$ to be maximally expressive about $Y$ while being maximally compressive about input data by:
@@
where $\beta$ is the Lagrange multiplier. It is obvious that Eq. 3 is the trade-off optimization between high mutual information and high compression Zhang et al. (2022a); Alemi et al. (2016). In our scenario, given a known small student model, the compression ratio is fixed. Therefore, we formulate the CoT distillation is an optimization problem as:
@@ 3.2 Variational Bounds of MI
According to  Poole et al. (2019); Covert et al. (2023), a tractable variational upper bound can be built by introducing a variational approximation $q(v)$ to the intractable marginal $p(v)$, thus:
@@
where $KL[\cdot||\cdot]$ denotes Kullback-Leibler divergence. The bound is tight when $q(v)=p(v)$. Then $KL(p(v)||q(v))=KL(p(v)||p(v))$, and it is a constant. Therefore, we can obtain the following inequality:
@@
Then we can write the MI as the following:
@@
Assuming that $p(v)$ is uniform distribution, $\sum p(v)p(z|v)\log p(v)$ is a constant. Then maximizing the $I(Z;V)$ can defined as:
$\begin{split}\max I(Z;V)&\varpropto\max\sum p(z|v)\log p(v|z)\\ &=\max(-\sum p(z|v)\log\frac{1}{p(v|z)})\\ &=\min(\sum p(z|v)\log\frac{1}{p(v|z)})\\ &=\min(\sum CE(z|v,v|z))\end{split}$ (9)
here $CE$ represents cross entropy. Therefore, based on this equation, we propose a new MI loss, which minimizes cross entropy to learn to maximize MI for CoT Distillation.
@@ 3.3 Training Loss
$f$ represents our proposed MI loss module, and $l$ denotes cross-entropy loss. As shown in Figure 1, the MI loss module module consists of softmax and max reduction layers. The softmax layer aims to generate an output distribution over representation learning space, while the max reduction aims to reduce the dimension of representation features of the label prediction task $\mathbb{R}^{m\times d}\rightarrow\mathbb{R}^{1\times d}$ and the dimension of representation features of the rationale generation task $\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{1\times d}$
@@ 4.1 Experimental Setting
We conducted extensive the experiments on 4 widely-used benchmark datasets across 3 different NLP tasks: e-SNLI Camburu et al. (2018) and ANLI Nie et al. (2020) for natural language inference; CQA Talmor et al. (2018) for commonsense question answering; SVAMP Patel et al. (2021) for arithmetic math word problems. We use rationale generated by PaLM 540B Chowdhery et al. (2023) collected and open-sourced by Hsieh et al. (2023) [^1] Data and DSS code are from https://github.com/google-research/distilling-step-by-step..
@@
Based on CoT properties and comparative experimental study in Hsieh et al. (2023), our work adopted T5-base (220 million) and T5-small (60 million) to the student models. $\alpha_{1}$ and $\alpha_{2}$ are set as $0.5$ and $\alpha_{3}$ is set as $0.1$. We trained our models on one A100 GPU with $80G$ memory. For T5 base model, the training time on full-size four dataset was about 14.4 hours. For T5 small model, the training times was from 8.6 hours.
@@
We compare our work with the state-of-the-art DSS Hsieh et al. (2023) by running its open-sourced code 1. We also compare with two most common methods in learning task-specific models: (1) Finetuning, which is the standard finetuning with the prevailing pretrain-then-finetune paradigm that finetunes a model with ground-truth labels via standard label supervision Howard and Ruder (2018). (2) Single-task, which the small model is distilled to predict labels with the teacher model’s predicted label. We run DSS, finetuning, and single-task with the same settings as Hsieh et al. (2023).
@@
Following the DSS work Hsieh et al. (2023), we adopt the accuracy as the performance metrics on all four datasets. A high accuracy indicates that the generated results are better. Besides it, we also adopt Expected Calibration Errors (ECE) and Average Confidence Scores to evaluate T5 calibration. Both higher scores indicate better. We adopt GPT-4 to evaluate Quality of CoT examples and subjective analysis. Please refer to our codes for more details.

### v3

@@ Abstract
Knowledge distillation, the technique of transferring knowledge from large, complex models to smaller ones, marks a pivotal step towards efficient AI deployment. Distilling Step-by-Step (DSS), a novel method utilizing chain-of-thought (CoT) distillation, has demonstrated promise by imbuing smaller models with the superior reasoning capabilities of their larger counterparts. In DSS, the distilled model acquires the ability to generate rationales and predict labels concurrently through a multi-task learning framework. However, DSS overlooks the intrinsic relationship between the two training tasks, leading to ineffective integration of CoT knowledge with the task of label prediction. To this end, we investigate the mutual relationship of the two tasks from Information Bottleneck perspective and formulate it as maximizing the mutual information of the representation features of the two tasks. We propose a variational approach to solve this optimization problem using a learning-based method. Our experimental results across four datasets demonstrate that our method outperforms the state-of-the-art DSS. Our findings offer insightful guidance for future research on language model distillation as well as applications involving CoT. Codes are available at https://github.com/xinchen9/cot_distillation_ACL2024.
@@ 3 Methodology
This section begins with an introduction to preliminaries of IB. Following this, we formulate our CoT distillation idea within the IB framework and propose a learning approach to optimize MI.
@@ 3.1 Preliminaries
Under the DSS framework, the prefixes [predict] and [explain] will be prepended to the input text, text, for tasks corresponding to label prediction and rationale generation, respectively. In the label prediction task, given the input [predict] + text along with predictive labels $\mathbf{Y}$, a representation feature $\mathbf{V}$, where $\mathbf{V}\in\mathbb{R}^{d}$, is trained using $\mathbf{Y}$. Similarly, in the rationale generation task, the input [explain] + text and rationale label $\mathbf{R}$ guide the training of a representation feature $\mathbf{Z}$, where $\mathbf{Z}\in\mathbb{R}^{d}$, using $\mathbf{R}$.
Our goal is to distill CoT knowledge from larger LLMs to smaller LLMs models. To achieve this, based on the basis of IB Tishby and Zaslavsky (2015); Zhang et al. (2022a); Wang et al. (2019), we model the DSS as following:
@@
where sampling observations $z\thicksim\mathbf{Z}$ and $v\thicksim\mathbf{V}$. Here $p(\cdot)$ represents the probability distribution.
To encourage CoT distillation to focus on the information represented in label $\mathbf{Y}$, we propose using IB to enforce an upper, bound $I_{c}$, on the information flow from the representation features $V$ to the representation features $Z$. This is achieved by maximizing the following objective:
@@
By employing Lagrangian objective, IB allows $Z$ to be maximally expressive about $Y$ while being maximally compressive regarding the input data, as follows:
@@
where $\beta$ is the Lagrange multiplier. Clearly, Eq. 3 demonstrates the trade-off optimization between high mutual information and high compression Zhang et al. (2022a); Alemi et al. (2016). In our scenario, given a predefined small student model, the compression ratio is fixed. Therefore, we formulate the CoT distillation as an optimization problem:
@@ 3.2 Variational Bounds of MI
According to  Poole et al. (2019); Covert et al. (2023), a tractable variational upper bound can be established by introducing a variational approximation $q(v)$ to replace the intractable marginal $p(v)$, demonstrated by:
@@
here $KL[\cdot||\cdot]$ denotes Kullback-Leibler divergence. The bound is tight when $q(v)=p(v)$. Consequently, $KL(p(v)||q(v))$ is equal to $KL(p(v)||p(v))$, which becomes zero. Therefore, we can derive at the following inequality:
@@
We can then express the MI from Eq. 5 as the follows:
@@
Assuming that $p(v)$ is uniform distribution for maximal entropy schroder2020estimating, the term $\sum p(v)p(z|v)\log p(v)$ is considered as a constant. This also allows for the omission of $p(v)$ in $\sum p(z|v)p(v)\log p(v|z)$. By combining Eq. 5 and Eq. 8, then maximization of $I(Z;V)$ can be expressed as:
$\begin{split}\max I(Z;V)&=\max\mathbb{E}_{p(z,v)}\Big{[}\log\frac{p(z|v)}{p(v)}\Big{]}\\ &\varpropto\max\sum p(z|v)\log p(v|z)\\ &=\max(-\sum p(z|v)\log\frac{1}{p(v|z)})\\ &=\min(\sum p(z|v)\log\frac{1}{p(v|z)})\\ &=\min(\sum CE(z|v,v|z))\end{split}$ (9)
here $CE$ represents cross entropy. Accordingly, CoT distillation in Eq. 4 is transformed into the problem outlined in the above equation. This problem can be addressed with a learning-based method. To tackle this issue, we have developed a new MI loss that minimizes cross-entropy of representation features of the rationale generation ($p(z|v)$) and representation features of the label prediction ($p(v|z)$), effectively maximizing MI during CoT distillation process.
@@ 3.3 Training Loss
$f$ represents our proposed mutual information (MI) loss module, and $l$ denotes the cross-entropy loss. As shown in Figure 1, the MI loss module consists of softmax and max reduction layers. The softmax function separately calculates the distributions for the outputs of the vocabulary spaces in the label prediction and rationale generation tasks. Subsequently, a max reduction operation is employed to reduce the dimensionality of the predicted outputs from both tasks to a uniform dimension for the loss calculation. Specifically, in the label prediction task, dimensions are reduced from $\mathbb{R}^{m\times d}$ to $\mathbb{R}^{1\times d}$, and in the rationale generation task, from $\mathbb{R}^{n\times d}$ to $\mathbb{R}^{1\times d}$.
@@ 4.1 Experimental Setting
We conducted the experiments on four widely-used benchmark datasets across three different NLP tasks: e-SNLI Camburu et al. (2018) and ANLI Nie et al. (2020) for natural language inference; CQA Talmor et al. (2018) for commonsense question answering; and SVAMP Patel et al. (2021) for arithmetic math word problems. We used rationale generated by PaLM 540B Chowdhery et al. (2023), which were collected and open-sourced by Hsieh et al. (2023) [^1] Data and DSS code are from https://github.com/google-research/distilling-step-by-step..
@@
Based on CoT properties and the comparative experimental study in Hsieh et al. (2023), our work adopted T5-base (220 million) and T5-small (60 million) to the student models. $\alpha_{1}$ and $\alpha_{2}$ were set as $0.5$ and $\alpha_{3}$ is set as $0.1$. We trained our models on one A100 GPU with $80G$ memory. For T5 base model, the training time for the full-size four dataset was approximately 14.4 hours. For T5 small model, the training times was approximately 8.6 hours.
@@
We compare our work with the state-of-the-art DSS Hsieh et al. (2023) by running their open-sourced code and include two other baseline reported in their work: (1) Standard Finetune, which involves using the prevailing pretrain-then-finetune paradigm to finetune a model with ground-truth labels through standard label supervision. Howard and Ruder (2018). (2) Single-task, which finetunes the model using both of the label and non-CoT rationale as supervision .
@@
Following the DSS work Hsieh et al. (2023), we adopt the accuracy as the performance metrics across all four datasets. Higher accuracy indicates better results. Besides accuracy, we also adopt Expected Calibration Errors (ECE) and Average Confidence Scores to evaluate calibration of the T5-base model. A lower ECE and higher Average Confidence Scores indicate better performance. We adopt GPT-4 to evaluate Quality of CoT examples and subjective analysis. Please refer to our codes for more details.